If money is worth 6% compounded bi-monthly, find the present value of Php125,700 for 2 years.

1
Expert's answer
2021-11-26T09:13:19-0500

compounded bi-monthly , n =6


"PV = \frac{FV}{(1+i)^n}"


"=\frac{Php125,700}{(1+0.06)^{6×2}}"


"=\frac{125,700}{(1.06)^{12}}"


"=Php62469.05"